OpenStudy (anonymous):

your car tire is rotating at 3.5 rev/s, when suddenly you press hard on the accelerator. after traveling 200 m, the tire's rotation has increased to to 6.0 rev/s. what was the tires angular acceleration?

OpenStudy (fibonaccichick666):

You should probably stick this in physics, but

OpenStudy (fibonaccichick666):

what is the formula for angular accel?

OpenStudy (anonymous):

dw/dt

OpenStudy (anonymous):

the problem is finding dt without the radius.

OpenStudy (fibonaccichick666):

and how do you calculate d(omega)?

OpenStudy (anonymous):

its 5pi but I can't figure out how long the process takes.

OpenStudy (fibonaccichick666):

all it wants is the accel though?>

OpenStudy (anonymous):

dw is not the acceleration, to find acceleration you need to know how long the change in omega took

OpenStudy (fibonaccichick666):

omega is ang velocity no? and average ang accel has multiple ways to calculate right? Assuming constant accel and given dw, and s

OpenStudy (fibonaccichick666):

dw=?

OpenStudy (fibonaccichick666):

ds/dt no?

OpenStudy (fibonaccichick666):

use that to calculate dt

OpenStudy (anonymous):

correct but ds is a function of r, which unknown

OpenStudy (praxer):

use the formula w^2-w_0^2=2 alpha theta where w is the final angular velocity, and w_0 is the initial angular velocity. Alpha is the angular acceleration and theta is the angular displacement. I am not sure ........will this work????.....:(

OpenStudy (anonymous):

ds = 200m. which makes d(theta) = 200/r and praxar that looked promising but I still end up with alpha*theta, neither of which I know. I need r to find theta

OpenStudy (praxer):

we will find the alpha so there is no need to have it I think so...:O

OpenStudy (praxer):

Ya but we need to get the radius... first...:)

OpenStudy (anonymous):

how do we get the radius?

OpenStudy (fibonaccichick666):

hmmm

OpenStudy (praxer):

@Dom89 Please check the question once again.. I think the radius must be given else it is just impossible to get hook of the solution..... i think so.......

OpenStudy (fibonaccichick666):

k try these, \(d\theta={d\omega}{dt}\)and

OpenStudy (praxer):

http://openstudy.com/study#/updates/51c57ae5e4b069eb00c573b6 Similar question but here the radius is given....

OpenStudy (fibonaccichick666):

\[\omega_f=\omega_i+2\alpha t\] and \[d\theta=\omega_i t+ .5 \alpha t^2\]

OpenStudy (fibonaccichick666):

3 variables, 3 eq

OpenStudy (anonymous):

don't you need one more equation than you have unknowns to find the unknowns?

OpenStudy (fibonaccichick666):

note \[\frac{d\theta}{dt}=d\omega\]

OpenStudy (fibonaccichick666):

no, just the same amount

OpenStudy (fibonaccichick666):

otherwise it's redundant

OpenStudy (fibonaccichick666):

and not LI

OpenStudy (anonymous):

sorry for the abrupt stop i got called to work. @praxer I checked the question, it definitely does not give the radius.

OpenStudy (anonymous):

thanks for the help guys but I asked a few people at uni and they all agreed it must be a mistake in the question. It's unsolvable without more info.

OpenStudy (fibonaccichick666):

Ok, that makes sense, otherwise you have a serious system of eq to solve haha